### v2

# Novel Strontium Carbides Under Compression

Abstract
Exploring the chemistry of materials at high pressures has lead to the discovery of previously unknown exotic compounds. Here, we systematically search for all thermodynamically stable Sr-C compounds under pressure (up to 100 GPa) using the ab initio evolutionary crystal structure prediction method. Our search lead to the discovery of hitherto unknown phases of SrC3, Sr2C5, Sr2C3, Sr2C, Sr3C2, and SrC. The newly discovered crystal structures feature a variety of different carbon environments ranging from isolated C anions and C-dimers to exotic polyatomic carbon anions including chains, stripes, and infinite ribbons consisting of pentagonal C5 and hexagonal C6 rings. Dynamical stability of all predicted compounds is confirmed by phonons calculations. Bader analysis unravels very diverse chemistry in these compounds and bonding patterns in some of them can be described using Zintl-Klemm rule.
pacs:
Valid PACS appear here
preprint: APS/123-QED
I Introduction
Chemical reactions that defy conventional chemical intuition can happen under extreme conditions, resulting in the emergence of rich phase diagrams and materials possessing intriguing properties, since pressure is known to greatly affect the chemistry of elements and cause the formation of exotic compounds McMillan (2002); Grochala et al. (2007). Well-known examples of such reactions involve the formation of sodium chlorides Zhang et al. (2013), a stable compound of helium and sodium Dong et al. (2017), platinum group metal nitrides Young et al. (2006), and superconducting metal hydrides Semenok et al. (2020). Nowadays, experimental high-pressure techniques allow one to explore materials chemistry up to terapascal pressure Dubrovinsky et al. (2022) and hence, confirm or deny theoretical predictions of exotic (from the chemical perspective) compounds. State-of-the-art computational techniques are reliable for predicting which compounds might be formed under certain thermodynamic conditions Oganov et al. (2019); Miao et al. (2020). For example, evolutionary crystal structure prediction has lead to the in silico discovery of transparent dense sodium, unexpected forms of sodium chloride, and various superconducting metal-hydrides, which were later confirmed experimentally. Ma et al. (2009); Zhang et al. (2013); Boeri et al. (2022)
Among all compounds previously investigated at high pressure, a particular focus was given to carbon-containing materials. Carbon (C) itself has the ability to form sp, sp2, and sp3-hybridized bonds and therefore exists in various allotropic forms, including known structures of graphite, diamond, graphene, fullerenes, nanotubes. Zhu et al. (2012); Rice et al. (1983); Hoffmann et al. (1991); Wang et al. (2015); Li et al. (2009); Boulfelfel et al. (2012); Wang et al. (2012); Hoffmann and Meyer (1992) High pressure alters the bonding patterns of carbides and leads to the formation of compounds with unusual carbon environments. Su et al. (2016); Krupka et al. (1969); Parasuk and Almlöf (1989) There are a large number of metal carbides containing acetylide group (C2 dimers) Toth (2014); Spedding et al. (1958); Sakai et al. (1981); Rundle et al. (1948); Akbar et al. (2024a); Feng et al. (2018), less common is the allylenide group (C3 unit). West et al. (1965); Mattausch et al. (1994); Fjellvaag and Karen (1992); Poettgen and Jeitschko (1991) For the binary systems Mg-C Strobel et al. (2014), Ca-C Li et al. (2015); Khandarkhaeva et al. (2024), Y-C Roszak and Balasubramanian (1996); Aslandukova et al. (2021); Gao et al. (2014), and La–C Su et al. (2016), an ab initio structure searches predicted the formation of unusual metal carbides with exotic C4, C5 units, C6 rings, and graphene carbon sheets.
Despite a thorough exploration of the chemical space of metal carbides in the past, recent experiments revealed several novel compounds in the Y-C and Ca-C systems Aslandukova et al. (2021); Khandarkhaeva et al. (2024). Motivated by these findings, we decided to construct a pressure compositions phase diagram of the Sr-C system. Compressed strontium shows unique structural and electronic properties Young Kim et al. (2012); Tsuppayakorn-Aek et al. (2015), but to the best of our knowledge, strontium carbides were not investigated under compression. ere, using a variable-composition structure prediction methodology the pressure-composition phase diagram of the Sr-C system was explored in order to fully understand the structural diversity and evolution of the C-C bonding types under high pressure in the Sr-C system. This resulted in several newly predicted stable stoichiometries (SrC3, Sr2C5, Sr2C3, Sr2C, Sr3C2, and SrC) with a diverse set of carbon anions: isolated C atoms in Sr2C, dimers in Sr3C2 and SrC, linear trimers in Sr2C3, chains in SrC, stripes with five-membered (C-pentagons) rings in SrC2. There are also infinite ribbons consisting of five-membered (C-pentagons) rings in SrC3, ribbons consisting of C-pentagons in SrC2, and ribbons made of C-pentagons and C-hexagons in Sr2C5. Here, we discuss structural and electronic properties of these compounds.
II Computational Methodology
Crystal structure prediction techniques allow one to determine pressure-dependent phase diagrams for a given chemical space and a set of thermodynamic conditions. Such methodology, called a variable-composition search, is not limited to the prediction of the ground state structure of a particular stoichiometry, but explores the whole chemical space of the system in a single calculation Oganov et al. (2011). Recently, by exploring pressure-composition phase diagrams of Cu-F, Ag-F, and Nd-F, we have shown that a variable-composition evolutionary crystal structure search allows one to find new stable compounds even after fixed-composition searches were done Rybin et al. (2021, 2022); Boeri et al. (2022).
Here, we used two crystal structure prediction codes, which both implement an idea of the evolutionary crystal structure prediction. One, is the field-leading crystal structure prediction package USPEX Glass et al. (2006); Oganov and Glass (2006); Lyakhov et al. (2013). Another one, is our own python-based implementation of the ab initio evolutionary algorithm, which has a functionality to perform variable-composition search. We call this implementation Sputnik (Structure prediction using theoretical kristallography) and details of the implementation will be published elsewhere, but we have to mention that workflow management in Sputnik is done using the FireWorks functionality Jain et al. (2015). The theoretical basis under the Sputnik’s implementation is identical to the USPEX’s one and it is presented in Deaven and Ho (1995); Oganov and Glass (2006); Oganov et al. (2011).
[…]
Once structure prediction with Sputnik and USPEX was done, we calculated the pressure-composition phase diagram with a resolution of 10 GPa, then we confirmed the dynamical stability of all novel structures with phonon calculations using the supercell approach and the finite displacement method Parlinski et al. (1997), as implemented in the Phonopy package Togo and Tanaka (2015). The electron localization function (ELF) was calculated for all the novel phases and visualized using VESTA.Momma and Izumi (2011) For aforementioned calculations we used VASP. Bader charge analyses was performed on total charge density obtained using VASP.Henkelman et al. (2006); Yu and Trinkle (2011); Sanville et al. (2007); Tang et al. (2009) The structural information for all discovered compounds, band structures, and phonon dispersion curves are presented in the Supplementary Materials (SM).
III Results
To determine stable phases in a variable-composition system, it is convenient to use the convex hull construction. Phases located on the thermodynamic convex hull are stable with respect to decomposition into elemental Sr and C or other Sr-C compounds. All the values of the chemical potentials, delimiting the fields of stability of compounds, are derived directly from the convex hull. Consequently, the convex hull construction is the cornerstone of the variable-composition structure prediction, which determines the stability of compounds at particular thermodynamic conditions Oganov et al. (2011).
Figure 1: Convex hull diagrams of the Sr-C system at 20 GPa (green disks) and 50 GPa (red squares). The enthalpy of formation $H$ is defined as: $\Delta H(Sr_{x}C_{y})=(H(Sr_{x}C_{y})-xH(Sr)-yH(C))/(x+y)$.
Figure 2: Pressure-composition phase diagram for the Sr-C system.
Convex hulls of the Sr-C system at 20 GPa and 50 GPa are presented in Fig. 1. The pressure-composition phase diagram of the Sr-C system is shown in Fig 2, which was obtained by calculating enthalpies of the most stable structures for each composition at a given pressure (steps of 5 GPa were used). According to this diagram, we found eleven hitherto unknown phases in the Sr-C system. These compounds have the following stoichiometries: SrC3, Sr2C5, SrC2, Sr2C3, SrC, Sr3C2, and Sr2C (for some stoichiometries there are a couple of different structures). These structures are shown in Fig. 3. All crystallographic parameters are given in SM. For all the newly predicted structures, calculated phonon dispersion relations confirmed their dynamical stability (see SM Fig.S1.).
The identified compounds have a variety of different carbon motifs, as demonstrated in Fig. 4. In the following, we consider the predicted phases in order of increasing carbon content. In order to analyze these structures, we recall that the length of the C - C bond depends on the bond order, and at 1 atm these lengths are 1.20 Å for the triple C–C bond, 1.33 Å for the double and 1.54 Å for the single C–C bond. Combining this knowledge with the results of Bader analysis, we unravel a very diverse chemistry.
Figure 3: Predicted crystal structures of stable Sr-C compounds. (a) $Pnma$ structure of Sr2C. (b, c) $C2/c$ and $P4/mbm$ structures of Sr3C2. (d, e) $C2/m$ and $P\bar{1}$ structures of SrC. (f, g) $C2/m$ and $P2_{1}/c$ structures of Sr2C3. (h, i) $Immm$ and $P\bar{1}$ structures of SrC2. (j) $P\bar{1}$ structure of Sr2C5. (k) $P2_{1}/m$ structures of SrC3. The green and brown spheres represent strontium and carbon atoms, respectively.
According to our calculations Sr2C crystallizes in the orthorhombic $Pnma$ space group. This structure is stable from 30 to 50 GPa, it is metallic, and contains isolated carbon atoms – the simplest carbon motif among all obtained compounds (see Fig. 3 (a)). Such stoichiometry was previously found in other metal-carbon compounds, for example, Ca2C and Mg2C. Li et al. (2015); Kurakevych et al. (2013) However, there is no reported theoretical or experimental data on Sr2C. This metallic phase (see SM Fig. S2 (a)) can be represented as a methanide with an idealized charge distribution of (Sr2+)2C4- and is in line with the Zintl concept.
Figure 4: Carbon patterns in the Sr-C system. (a) The isolated carbon anions in the Pnma structure of Sr2C. (b) Carbon dimers (dumbbells) observed in the C2/c structure of Sr3C2 and C2/m structure of SrC. (c) The carbon trimer occurs in the C2/m structure of Sr2C3. (d) Zigzag C6 groups observed in $P\bar{1}$ structure of SrC at 50 GPa. (e) Carbon stripes in the P-1 structure of SrC2 at 50 GPa. (f) Carbon ribbons with hexagonal C-rings in the Immm structure of SrC2 at 90 GPa. (g) Carbon ribbons with pentagonal C-environment in the Pnma structure of SrC3 at 20 GPa. (h) Carbon ribbons with pentagonal and hexagonal C-rings in the $P\bar{1}$ structure of Sr2C5 at 50 GPa.
Sr3C2 crystallizes in the tetragonal $P4/mbm$ space group at 30 GPa and transforms into a structure with a monoclinic $C2/c$ space group at 50 GPa. The structure of the newly identified Sr3C2 is of the U3Si2 type, a common arrangement found in silicides Zachariasen (1948), borides Riabov et al. (1999), and intermetallic compounds Chai and Corbett (2011). Both structures are metallic and contain carbon dimers as presented in Figs. 3 (b, c). Carbides of calcium with this stoichiometry and space groups have been predicted previously Li et al. (2015). In $P4/mbm$-Sr3C2 phase, the structure contains doubly bonded C2 groups (the C-C distance is 1.39 Å at 30 GPa) carrying an ideal charge of -4, receiving four electrons from Sr atoms. This leaves two electrons per C atom to establish C–C double bonds in the metallic framework. Hence, Sr3C2 (as well as Ca3C2) can be called an ethanide - at normal conditions ethanides are unknown unlike methanides and acetylides. A carbon atom with two additional electrons becomes isoelectronic with an oxygen atom, similar to the configuration found in the doubly bonded O2 molecule. The Sr atoms are located in the same plane as the C2 dumbbells, collectively forming the Cairo pentagonal tiling, which consists of (Sr)2C3 pentagons (see Fig. 5 (a)). As shown in Fig. 5 (a), Sr atoms are four-fold coordinated by C atoms, with a Sr-C distance of 2.63 Å at 30 GPa. The electron localization function indicates strong covalent bonding between the carbon atoms in the dimers and an ionic Sr-C interaction (see Fig. 5 (c)). Above 50 GPa the $P4/mbm$-Sr3C2 transforms into $C2/c$-Sr3C2 and the Cairo pentagonal tiling gets distorted. This distortion is accompanied by an increase in the C-C bond length of the C-dimers to 1.46 Å. Electron density maps in the plane containing the C2 dimers reveal slightly higher ELF values between carbon atoms in the dimers of the $P4/mbm$-Sr2C3 phase compared to the $C2/c$-Sr2C3 phase (See SM Fig. S6). This suggests a higher bond order in the $P4/mbm$-Sr2C3 dimer.
Metallic SrC exhibits two thermodynamically stable phases below 100 GPa. The low pressure monoclinic $C2/m$ phase on further compression transitions into a low-symmetry triclinic $P\bar{1}$-SrC structure. At 30 GPa the structural formula of $C2/m$ is Sr2(C2), featuring a doubly bonded C2 group with a C–C bond length of 1.34 Å. This C2 group carries an ideal charge of -4 (Bader charge -2.175), perfectly balancing the charge of two strontium atoms, and $C2/m$-SrC is classified as an ethylenide. The $P\bar{1}$-SrC structure is particularly intriguing due to its novel zigzag C-chain with C–C bond lengths ranging from 1.45 to 1.50 Å, implying bond orders between 1 and 2. The bond angles are 125.6 and 121.3, which are remarkably close to 120 expected for sp2-hybridized carbon. The ELF maps clearly indicate strong covalent bonding between the carbon atoms and ionic bonding between Sr and C atoms (see SM Fig.S5).
Figure 5: Crystal structure of $P4/mbm$-Sr3C2 at 30 GPa is depicted along with the computed electron localization function (ELF). Green and brown spheres represent strontium and carbon atoms, respectively. (a) Sr-C plane is a the Cairo pentagonal tiling formed by Sr and C atoms. (b) interatomic distances within the Sr-C plane. (c) 2D ELF of the $P4/mbm$-Sr3C2 structure.
At lower pressures of about 10 GPa Sr2C3 becomes stable and crystallizes in the monoclinic $C2/m$ space group, and its structure contains allylenide-groups (C3 units). Beyond 20 GPa, it transforms to the structure with $P2_{1}/c$ space groups, which is stable up to 50 GPa. These structures are the only semiconducting compounds found in this study (see SM Fig.S2). Usually, C3 trimers are linear [C=C=C]4- groups with a C-C bond length of 1.34-1.35 Å (e.g., in Mg2C3), which is close to that in gaseous allene (1.335 Å) Almenningen et al. (1959). The structure of $C2/m$-Sr2C3 at 20 GPa consists of Sr2 layers connected by nearly linear, symmetric C3 groups featuring double C-C bonds, with C-C distances measuring 1.31 Å. In this configuration, the ideal charge of the C3 group is -4 (Bader charge -2.625), which precisely balances the total charge of the two Sr atoms in the formula, with one Sr atom having +1.314 Bader charge. The central carbon atom in the C3 group should be neutral and the terminal carbons should have -2 charges as in gaseous allene. However, the central atom of the C3 unit in Sr2C3 exhibits higher negative Bader charge of -1.225, while the terminal carbon atoms show lower Bader charges of -0.700 (see SM Tab. 1). This discrepancy is attributed to the significant ionic interactions between the Sr atom and the central carbon atom of the C3 unit. This can be seen in the calculated 2D ELF of these structures (see SM Fig. S3).
SrC2 becomes stable at 30 GPa and crystallizes in the low-symmetry $P\bar{1}$ space group. This structure contains exotic carbon stripes with fused five-membered C rings – the structural motif is presented in Fig. 4 (e). In this structure, two layers of Sr atoms are stacked between two layers of these stripes. The C - C distances within the C rings range from 1.41 to 1.48 Å at 30 GPa, indicating the bonds intermediate between single and double. Upon compression, at 90 GPa there is a phase transition $P\bar{1}$-SrC2 $\rightarrow$ $Immm$-SrC2. The $Immm$-SrC2 structure contains extended carbon ribbons consisting of fused hexagonal C rings as presented in Fig. 4(f). These carbon ribbons are planar, with C–C distances of 1.43 Å and 1.44 Å at 90 GPa. The $Immm$-SrC2 phase shares the same structure as the previously observed HP-CaC2 ($Immm$Li et al. (2015) and the predicted high-pressure YC2 ($Immm$)  Feng et al. (2018). In the low symmetry $P\bar{1}$-SrC2 phase, the strontium atom is surrounded by 12 carbon atoms forming a heptagon and pentagon above and below the Sr atom (see Fig. 6 (a)). In the $Immm$-SrC2 phase, the strontium atoms have 12 carbon neighbours arranged in a hexagonal prism with unequal bases (see Fig. 6 (c)). As was mentioned above, these C atoms are organized into poly-cyclic units of hexagonal rings, forming planar ribbons and as a result, the carbon atoms create exotic, one-dimensional extended poly-anions. The hexagon features two angles of 122.5 and four angles of 118.8. These angles, being close to 120, along with the similar C–C distances, suggest that the carbon atoms are sp2 hybridized. The shared edges of hexagons in $Immm$-SrC2 are slightly shorter (by approximately 0.01 Å) than the non-shared edges. This observation aligns with calculations in previously predicted isostructural high-pressure carbides (such as DyC2, CaC2, and YC2), which show shorter shared edges by about 0.02–0.03 Å  Akbar et al. (2024b). Bader charge analysis (See SM Tab. 1) suggests charge transfer from cationic Sr to C atoms. ELF maps of carbon ribbons in (see Fig. 6 (b, d) ) suggest strong covalent bonds between carbon atoms and show localization of electrons on C-C bonds in pentagonal ribbons and hexagons.
Figure 6: The crystal structure of SrC2 depicted along with the computed Electron Localization Function (ELF). Green and brown spheres represent strontium and carbon atoms, respectively. (a) Coordination of Sr atoms by C atoms, involving a pentagonal base, (b) 2D electron localization function (ELF) containing the pentagonal ribbons, (c) coordination environment of Sr atom in Immm phase of SrC2 (d) ELF containing the carbon ribbons in the $Immm$-SrC2 phase.
Sr2C5 crystallizes in the low-symmetry space group $P\bar{1}$. C atoms in this structure feature extended ribbons consisting of C-pentagons and C-hexagons, as shown in Fig. 4(g, h).The C-C bond lengths range from 1.39 Å to 1.44 Å, indicating an intermediate bond order between single and double C-C bonds. In the Sr2C5 structure, one Sr atom has +1.248e Bader charge and is 13-fold coordinated, while another Sr atom has +1.202e Bader charge and is 12-fold coordinated (see Fig. 7(a)). In the fused C-hexagon and C-pentagon, carbon atoms are negatively charged. Therefore, it is evident that there is charge transfer from Sr to C atoms. Sr atom with 13-fold coordination is surrounded by the C-hexagon and a heptagon, while the Sr atom with 12-fold coordination is surrounded by the C-pentagon and a heptagon. In the C-hexagons, the four inner C atoms have three C-C covalent bonds each, while the two outer C atoms have two C-C covalent bonds and one lone pair of electrons each. The C-C bonds in the C-hexagon form two different angles $\alpha$ and $\beta$, where $\alpha$=112 and $\beta$= 124. In C-pentagons, the inner four C atoms have three C-C covalent bonds, while the outer C atom has two C-C covalent bonds and a lone pair of electrons. The lone pairs of electrons can be seen in the 2-D electron localization function (see Fig. 7 (b)).
Figure 7: (a) Coordination environment of the Sr atoms with the fused C-hexagons and C-pentagons in $P\bar{1}$-Sr2C5 (b) 2D ELF of the fused extended C-ribbons, consisting of C-hexagons and C-pentagons
Finally, SrC3 crystallizes in the $Pnma$ space group. In this structure C atoms form infinite ribbons consisting of pentagonal carbon rings (see SM Fig. S4 (a)). Bader analysis indicates a charge transfer from strontium to carbon atoms, with strontium atoms acquiring a positive Bader charge of +1.343 (see SM Tab. 1). The carbon atoms, in contrast, have negative charges, with those not involved in ring fusion showing higher electron concentration. This is also evident in the 2D ELF plot (see SM Fig. S4 (b)).
In light of the appreciable diversity of the observed carbon motifs (see Figure 4), it is instructive to check if at least some of these configurations can be explained using conventional chemical wisdom. For that purpose, we investigated the applicability of the Zintl-Klemm (ZK) concept, originally conceived to rationalize bonding patterns of polyanions, to the newly identified strontium carbides. To assess the validity of the ZK rules, the average charge of a C-anion will be determined by assigning a formal $+2$ charge to the Sr atoms (Sr2+) and imposing charge neutrality of the formal SrxCy unit. For the example of Sr2C this would lead to a formal charge of the carbon anion of $-4$. By subsequently equating the structural motifs formed by that C-anion to the isoelectronic, neutral species, in the case of C4- that would be a neon atom, the structure of the polyanion is inferred. Since neon is a noble gas, and Ne-Ne bonds cannot be formed, we expect no C-C bonds in Sr2C, which we indeed observe. We performed this analysis for all herein discussed stoichiometries, which we have summarized in Table 1. We find that for some of the strontium carbide compounds, namely Sr2C, Sr3C2 and SrC, the C-anion structure is consistent with the motifs predicted from the ZK rule. This finding is also in agreement with a previous investigation of high-pressure Ca-C phases Li et al. (2015). In the other cases, however, we find that the ZK concept cannot be directly applied, since for some stoichiometries including Sr2C3, SrC3 and Sr2C5, the average formal charge on a C-atom would be fractional. Double and triple bonds also lead to exceptions from the ZK rule. A particularly noteworthy example is the case of SrC2, in which the carbon species would be isoelectronic to a nitrogen atom. As a consequence one would expect each carbon atom to form three covalent bonds. As illustrated in Figure 4, we find that the carbon polyanion in $Immm$-SrC2 consists of planar, napthalene-like arrangements. The planarity and the similarity of all C-C bond lengths in that polyanion suggests that that configuration is aromatic. This implies that most of the carbon atoms form one single and one double bond, so that the bonding of these atoms in $Immm$-SrC2 is consistent with the ZK prediction. The two carbon atoms shared between the two rings, however, must then consequently have two single bonds and one double bond, which is not compatible with the isoelectronic N-atom. The same argument applies to the planar structure of the anion in the $P\overline{1}$ phase of SrC2, so that both SrC2 phases are in contradiction with the ZK rule.
Compound Formal C-charge isoelectronic to Expected anion structure ZK rule fulfilled?
Sr2C $-4$ Ne mono-atomic Yes
Sr3C2 $-3$ F dimer Yes
SrC $-2$ O dimer Yes
Sr2C3 $-4/3$ N/A N/A N/A
SrC2 $-1$ N trigonal non-planar No
SrC3 $-2/3$ N/A N/A N/A
Sr2C5 $-2/5$ N/A N/A N/A
Table 1: Applicability of Zintl-Klemm (ZK) rules to the newly identified stoichiometries of strontium carbide. In cases where the formal charge was found to be fractional, the Zintl-Klemm rules cannot be applied (”N/A”) in a straightforward manner.
IV Conclusion
In summary, we have produced the first complete pressure-composition phase diagram for Sr-C compounds at pressures up to 100 GPa. As a result we predicted eleven novel thermodynamically stable phases in the Sr-C system. Using electronic structure calculations and by performing Bader charge analysis we unraveled very diverse chemistry in these compounds. Carbon atoms in these compounds form a diverse set of different environments: isolated C atoms in Sr2C (methanide), dimers in Sr3C2 (ethanide) and SrC (ethylenide), linear trimers in Sr2C3 (allylenide), chains in SrC, stripes with five-membered (C-pentagons) rings in SrC2. There are also infinite ribbons consisting of five-membered (C-pentagons) rings in SrC3, ribbons consisting of C-pentagons in SrC2, and ribbons made of C-pentagons and C-hexagons in Sr2C5. Bonding patterns in some of the obtained compounds can be described using the Zintl-Klemm rule.
We note that while powerful computational methods, such as evolutionary crystal structure prediction used in this work, are capable of reliably predicting these exotic Sr-C compounds under pressure, experimental validation of the obtained results should be done.
V Acknowledgements
N.R. and A.S. acknowledge funding from the Russian Science Foundation (Project No. 23-13-00332). A.R.O. and P.G. acknowledge funding from the Russian Science Foundation (Project No. 19-72-30043). E.M. thanks the NOMAD Laboratory at the FHI of the Max Planck Society for support of the project.